### v1

# Extracting and Transferring Abilities For Building Multi-lingual Ability-enhanced Large Language Models

Abstract
Multi-lingual ability transfer has become increasingly important for the broad application of large language models (LLMs). Existing work highly relies on training with the multi-lingual ability-related data, which may be not available for low-resource languages. To solve it, we propose a Multi-lingual Ability Extraction and Transfer approach, named as MAET. Our key idea is to decompose and extract language-agnostic ability-related weights from LLMs, and transfer them across different languages by simple addition and subtraction operations without training. Specially, our MAET consists of the extraction and transfer stages. In the extraction stage, we firstly locate key neurons that are highly related to specific abilities, and then employ them to extract the transferable ability-specific weights. In the transfer stage, we further select the ability-related parameter tensors, and design the merging strategy based on the linguistic and ability specific weights, to build the multi-lingual ability-enhanced LLM. To demonstrate the effectiveness of our proposed approach, we conduct extensive experiments on mathematical and scientific tasks in both high-resource lingual and low-resource lingual scenarios. Experiment results have shown that MAET can effectively and efficiently extract and transfer the advanced abilities, and outperform training-based baselines methods. Our code and data are available at https://github.com/RUCAIBox/MAET.
1 Introduction
Large language models (LLMs) have recently shown remarkable performance on various general tasks, e.g., text generation and question answering (Zhao et al., 2023; OpenAI, 2023; Dubey et al., 2024). Despite the success, LLMs are still struggling to solve complex tasks (e.g., mathematical reasoning), which require LLMs to possess specific advanced abilities (e.g., deductive reasoning) and knowledge (e.g., mathematical theory) (Yue et al., 2024; Lu et al., 2022). To address it and further improve LLMs, existing work either collects the related data to train LLMs (Du et al., 2024; Chen et al., 2024a), or merges the parameters of existing well-performed LLMs to transfer their advanced abilities into one single model (Ilharco et al., 2023; Yadav et al., 2023; Yu et al., 2024a).
Despite the success, it is not easy to collect sufficient training corpus or well-trained LLMs related to specific abilities, especially in multi-lingual scenarios. Especially, some popular languages (e.g., English) have dominated the linguistic expressions of the open web data, and the amount of available data for low-resource languages (e.g., Bengali or Telugu) is highly limited (Patzelt, 2024; Mirashi et al., 2024). Fortunately, existing work (Zhao et al., 2024; Schäfer et al., 2024) has revealed that the learned knowledge from one language by LLMs could be inherited and leveraged by other languages. For example, Llama-series LLMs are trained mainly on English texts, while they can also solve the tasks based on other languages. Such a finding has been widely explored in either improving the overall performance of multi-lingual LLMs (Schäfer et al., 2024) or enhancing fine-grained knowledge (Chen et al., 2024a). However, the related work mostly relies on training with ability-related corpus in the target language, which is not always available for low-resource languages.
To conduct a more effective ability transfer, our idea is to learn and extract the “ability-specific weights” that preserves the knowledge about specific abilities for the LLM. If such ability-specific and language-specific weights could be decomposed, it is achievable to transfer the required abilities into target languages by just combining the corresponding weights, even building a multi-lingual ability-enhanced LLM like building blocks. Based on this idea, in this paper, we propose a Multi-lingual Ability Extraction and Transfer approach, named as MAET. Specifically, our approach consists of two major stages, i.e., ability extraction and transfer stage. In the extraction stage, we first locate the abilities-related neurons and leverage related corpus in a reference language to continually pre-train the LLM on these identified neurons. Then, based on the LLM trained on the general corpus, we devise the formula to extract the ability-specific weights. In the transfer stage, we utilize the ability-related weights to select related parameter tensors, and design a specific model merging strategy by interpolating linguistic and ability-specific weights. In our approach, we only need ability-specific corpus from any rich-resource language and general multi-lingual corpus, which can effectively mitigate the data scarcity issues in low-resource languages.
To assess the effectiveness of our approach, we conduct the evaluation based on two comprehensive reasoning benchmarks, namely Multi-lingual Grade School Math (MGSM) (Shi et al., 2023) and science tasks from multi-lingual MMLU (Lai et al., 2023) as the evaluation benchmarks. According to the evaluation results, the proposed approach MAET outperforms other competitive baseline methods (e.g., continual pre-training (Gururangan et al., 2020) and model merging methods with task vectors (Ilharco et al., 2023), achieving the 9.1% relative improvement upon the backbone LLM. In addition, our approach can work well with relatively fewer training data, demonstrating an improved efficiency in practice. In conclusion, our contribution can be summarized as follows,
(1) Our research has found that advanced abilities can be extracted from the single-lingual corpus and transferred across languages without the multi-lingual ability-related corpus, enabling to efficiently empower LLMs with special advanced abilities.
(2) We propose an effective and efficient approach named MAET, which first identifies and extracts the ability-related weights in LLMs and then only leverages simple addition and subtraction operations to build a multi-lingual ability-enhanced LLM.
(3) Our approach MAET achieves better performance than the competitive baseline methods (e.g., continual pre-training and model merging with task vector) in multi-lingual complex reasoning tasks, including mathematical reasoning tasks and scientific reasoning tasks.
2 Related Work
We introduce the related work from the following three perspectives:
Continual Pre-training. Although LLMs have shown remarkable performance on various downstream work, they still struggle in several specific tasks, e.g., complex reasoning tasks (Paster et al., 2024; Shao et al., 2024) or low-resource lingual scenarios (Hedderich et al., 2021; Panchbhai & Pankanti, 2021). To adapt LLMs pre-trained on the general corpus to multi-lingual scenarios or specific tasks, existing work (Luo et al., 2022; Taylor et al., 2022; Zhao et al., 2022; Zhang et al., 2024a) has collected the corresponding corpus to continually pre-train (CPT) LLMs. During the continual pre-training process, the mixture strategy between the general corpus and the ability-related corpus should be carefully considered to prevent hurting the general abilities of LLMs (Ye et al., 2024; Xie et al., 2023; Chen et al., 2024a; Siriwardhana et al., 2024). However, previous study Chang et al. (2024); Lu et al. (2023) has pointed out that it is difficult to collect the required corpus, especially for low-resource language corpus. Therefore, synthesizing data from powerful LLMs is widely utilized to expand the task-specific training corpus (Chen et al., 2021b; Yu et al., 2024b; Zhou et al., 2024a). Besides, because of the limitation of computation resources, a series of approaches (Hu et al., 2022; Li & Liang, 2021; Dettmers et al., 2023) only train several parameters to reduce the expenses. In this work, we focus on adapting LLMs to multilingual complex reasoning scenarios through continually pre-training LLMs on the single-lingual task-specific corpus.
Knowledge Editing. According to lottery ticket hypothesis (Frankle & Carbin, 2019), training a small number of model parameters will achieve comparable or even better performance on downstream tasks. Existing study (Du et al., 2024; Wang et al., 2024b; Gong et al., 2024) has leveraged the inner information of LLMs, e.g., gradient or cosine similarity between different hidden states, to select and train the related sub-network. Besides, the probe (i.e., a newly initialized parameter) can be implemented to detect the knowledge of LLMs and process targeted repair (Wang et al., 2024a; Jiang et al., 2024). However, Since these approaches need to calculate and select a sub-network of each training instance, which might cause the instability of the training process, several study (Chen et al., 2024b; Zhang et al., 2024b) pointed out that the task-related sub-network can be determined before the training process, and only updating the value of the corresponding neurons can achieve better performance. In this work, we focus on editing the task-specific neurons of LLMs to improve the corresponding capacities in multi-lingual scenarios.
Model Merging. Given that the CPT process will bring huge computational expenses, previous work leveraged model merging techniques to integrate different abilities (e.g., mathematical reasoning and code synthesizing) into one model (Yang et al., 2024; Xu et al., 2024; Stoica et al., 2024). During the merging process, the interference between different LLMs might be conflict with each other and affect the final performance. Therefore, researchers proposed the clip (Yadav et al., 2023) or randomly dropout (Yu et al., 2024a) mechanism to mitigate the performance decrease. Moreover, the selection of the hyper-parameters (e.g., weight of each model) is the challenge of the model merging process, and previous work (Zhou et al., 2024b; Matena & Raffel, 2022) utilized the inner parameters of LLMs or external matrixes to determine the hyper-parameters. Inspired by the above work, we try to locate the task-related sub-networks of LLMs and transfer the advanced abilities.
3 Preliminary
Despite that LLMs exhibit excellent performance on general tasks, they still have limited advanced abilities, e.g., mathematical and scientific reasoning abilities. A typical approach to enhancing these abilities is to continually pre-train (CPT) LLMs with ability-related corpus. However, such training data might not always be available or sufficient, especially for minor domains (e.g., Bengali). In this work, we focus on the task of ability extraction and transfer by continual pre-training and merging LLMs. Concretely, LLMs are trained on collected corpus from a certain domain, and we aim to only transfer its learned advanced capabilities to another target domain (Zhuang et al., 2021; Farahani et al., 2021) without further training efforts. In this work, we mainly study the cross-lingual scene where the linguistic-agnostic advanced ability and linguistic abilities should be extracted and transferred, to help build a unified multi-lingual ability-enhanced LLM.
Formally, for a certain ability $A_{i}$ and a set of languages $L=\{L_{0},L_{1},\dots,L_{n}\}$, we assume that the general corpus of all languages can be collected, denoted as $\mathcal{C}_{\text{general}}=\{\mathcal{C}_{L_{0}},\mathcal{C}_{L_{1}},\dots,\mathcal{C}_{L_{n}}\}$, while the ability-related corpus is only available in language $L_{0}$ (i.e., English), denoted as $\mathcal{C}_{L_{0},A_{i}}$. Based on the above corpora, our goal is to extract and transfer the advanced ability $A_{i}$ from language $L_{0}$ and linguistic abilities from other languages $L_{1},\dots,L_{n}$, into a unified LLM.
4 Approach
Figure 1: The framework of our approach MAET, including extraction and transfer stages. In the extraction stage, we first locate the key neurons, and utilize the single-lingual ability-related corpus and general corpus to train the LLM on these neurons to obtain the ability-specific weight. Then, we remove the parameter tensors related to language knowledge in the ability weight and transfer the remaining to the base LLM. After these stages, we can obtain a powerful LLM with advanced abilities that can solve the corresponding tasks in multi-lingual scenarios.
In this section, we propose the Multi-lingual Ability Extraction and Transfer approach, named as MAET, which can effectively transfer the advanced abilities from single-lingual LLMs, to build a multi-lingual ability-enhanced LLM. The key motivation of our approach is to identify and extract ability-related neurons or weights, and transfer the target abilities into a LLM in an efficient way. The overall framework of MAET is presented in Figure 1.
4.1 Ability-related Weights Extraction
In this part, we aim to locate and learn ability-related parameter weights within an LLM, to enable efficient transferring of the ability into other LLMs. Concretely, it consists of two major steps, i.e., key neurons locating and ability-related parameter weights learning, which are detailed in the following.
Locating the Key Neurons. The gradient of each neuron in LLMs can be utilized to estimate its correlation degree with specific task ability , we select those with high gradient values as key neurons. To this end, we first use a small amount of ability-related corpus $\mathcal{C}_{L_{0},A_{i}}^{\prime}$ to continually pre-train the LLM. During training, the LLM learns the language modeling task and each neuron is updated by the gradients associated by the training instances. Due to the high cost of calculating the accumulation of gradient at each training step, we calculate the value changes of the LLM neurons before and after the training process to approximate the importance. Formally, the importance function $I(A_{i},\theta_{j})$ of neurons can be computed as:
$I(A_{i},\theta_{j})=\sum_{d_{k}\in\mathcal{C}_{L_{0},A_{i}}^{\prime}}{\text{Gradient}\left(\theta_{j},d_{k}\right)}\approx\lambda\cdot\parallel\tilde{\theta}_{j}-\theta_{j}\parallel,$ (1)
where $d_{k}$ denotes the $k$-th instance of training corpus $\mathcal{C}_{L_{0},A_{i}}^{\prime}$ and $\tilde{\theta}_{j}$ denote the value of the $j$-th neuron of LLM after training, respectively. Based on it, we rank all neurons according to their importance scores, and then select the top $k_{1}\%$ ones into the set $\mathcal{N}_{A_{i}}$ as the key neurons.
Learning Ability-specific Weights. Based on the identified key neurons in $\mathcal{N}_{A_{i}}$, we further learn the ability-specific parameter weights. Our motivation is to decompose the parameter weights according to their changes before and after the LLM has mastered a specific ability, which is achievable owing to the modularity and composition nature of the LLM parameter matrices (Yu et al., 2024a; Shazeer et al., 2017). First, we utilize the key neurons locating method mentioned above to extract the ability-related neuron set $\mathcal{N}_{A_{i}}$, and also obtain the language-related neuron set $\mathcal{N}_{L_{0}}$ via the same way. Then, we train the LLM with the mixture of ability-related corpus and general corpus on the key neuron set $\mathcal{N}_{A_{i}}\bigcup\mathcal{N}_{L_{0}}$ and $\mathcal{N}_{L_{0}}$ respectively, to obtain two specific models, denoted as $\text{LLM}_{A_{i},L_{0}}$ with parameters $\Theta_{A_{i},L_{0}}$ and $\text{LLM}_{L_{0}}$ with parameters $\Theta_{L_{0}}$. Next, we measure the parameter changes between the backbone and the trained models, and obtain the ability-specific weights via the parameter decomposition operation as:
$R(A_{i})=\alpha\cdot\underbrace{(\Theta_{A_{i},L_{0}}-\Theta_{o})}_{\text{Ability \& language difference}}-\beta\cdot\underbrace{(\Theta_{L_{0}}-\Theta_{o})}_{\text{Language difference}},$ (2)
where $\alpha$ and $\beta$ are tunable coefficients to balance the two parts of weight differences, and $\Theta_{o}$ denote the original parameters of the LLM, which serves as the reference for parameter decomposition. As we only train the parameters within the neuron set, its weight difference should preserve the knowledge about the corresponding ability. Thus, it can be regarded as the ability-specific parameter representations, and is promising to transfer the ability into other LLMs by the addition operation.
1
Input : Single-lingual ability-related corpus $\mathcal{C}_{L_{0},A_{i}}$, multi-lingual general corpus $\mathcal{C}_{L_{0}},\mathcal{C}_{L_{1}},\dots,\mathcal{C}_{L_{n}}$, and the parameters of the backbone model $\Theta_{o}$.
Output : A well-trained multi-lingual ability-enhanced LLM.
2
// Ability-related Weights Extraction
3 $\theta^{\prime}\leftarrow\text{CPT}(\mathcal{C}_{L_{0},A_{i}},\Theta_{o}$);
4
5for $j$-th neuron in $\Theta_{o}$ do
6       Calculate the importance score of the corresponding neuron using Eq. 1;
7
8
9Identify the key neuron set $\mathcal{N}_{A_{i}}$;
10
[…]
12
[…]
14
15Learning the ability-specific weight $R(A_{i})$ using Eq. 2;
16
// Multi-lingual Ability Transfer
17
18Obtaining the multi-lingual weight $R_{Lang}$ using Eq. 3;
19
20for $j$-th parameter tensor in LLM do
21       Calculate the correlation using Eq. 4;
22
23
24Identify the ability-related parameters $\mathcal{T}$;
25
26Transfer the ability to multi-lingual scenarios using Eq. 5;
27
28Obtain the well-trained multi-lingual ability-enhanced LLM.
Algorithm 1 The complete procedure of our proposed approach.
4.2 Multi-lingual Ability Transfer
After obtaining the ability-related weights, in this part, we utilize them to transfer and integrate the abilities, to build a multi-lingual ability-enhanced LLM.
Ability-related Parameter Tensor Selection. Although we can locate the ability-related key neurons, it is still hard to avoid the involvement of irrelevant ones. Our empirical studies have found that neuron-level features are easy to be affected by the noisy data. Therefore, we consider identifying ability-related parameter tensors, which correspond to the parameter matrices within the LLM. Specifically, we firstly leverage the ability-related weights of languages $R(L_{1}),\dots,R(L_{n})$ to obtain the multi-lingual weight $R_{Lang}$. Given that large models have varying levels of proficiency in different languages, we use the hyper-parameters $\mu_{1},\dots,\mu_{n}$ to tune this process as:
$R_{Lang}=\sum_{i=1}^{n}{\mu_{i}\cdot R(L_{i})},$ (3)
where $R(L_{i})$ preserves the linguistic ability of language $L_{i}$ learned based on Equation 2. Therefore, $R_{Lang}$ can be considered as the general language ability of LLMs that spans multiple languages. As we aim to find he parameter tensors that have low linguistic effects but focus on the desired abilities (e.g., mathematical reasoning), we rank all the tensors according to their similarities with $R_{Lang}$, and pick up the last $k_{2}\%$ ones. Formally, for tensor $\tau_{i}$, we calculate the cosine similarity of this parameter between $R(A_{i})$ and $R_{Lang}$, as follows,
$S(\tau_{i})=\texttt{sim}\left(R(A_{i})[\tau_{i}],R_{Lang}[\tau_{i}]\right),$ (4)
where we use the dot product to implement the similarity function $\texttt{sim}\left(\cdot\right)$. After obtaining the similarity of all tensors, we rank them in a descending order based on the similarity values, and then select the last $k_{2}\%$ parameters into the set $\mathcal{T}$ as the ability-related parameters.
Building Multi-lingual Ability-enhanced LLM. Based on the selected ability-related tensors $\mathcal{T}$, we design the model merging process by interpolating ability weights and multi-lingual weights, to build the multi-lingual ability-enhanced LLM. Formally, the final parameter tensors of the target LLM are computed as:
$\tilde{\tau}_{i}=\tau_{i}^{(o)}+\begin{cases}\gamma\cdot R(A_{i})[\tau_{i}]+\eta\cdot R_{Lang}[\tau_{i}],&\tau_{i}\in\mathcal{T}\\ R_{Lang}[\tau_{i}],&\tau_{i}\notin\mathcal{T}\end{cases},$ (5)
where $\tau_{i}^{(o)}$ denotes the original value of parameter tensor $\tau_{i}$, and $\gamma$ and $\eta$ are tunable hyper-parameters. This formula can be explained in two different cases. When a parameter tensor serves as the major role for specific abilities, we update it by adding both ability- and linguistic-related weights; otherwise, we simply enhance it with multi-lingual weights. In this way, we can derive a more powerful LLM that is equipped with the multi-lingual abilities and specific advanced abilities.
Aspects CPT MoE LoRA MoL TV MAET (Ours)
MLAR Corpus Yes Yes Yes Yes Yes No
Tuning Parameters Full Full Low-Rank Low-Rank Full Ability-related
Ability Composition No No No No Yes Yes
Ability Transfer No No No No No Yes
Table 1: The difference between our proposed MAET and the methods in previous work (i.e., CPT (Hu et al., 2022), Mixture-of-Expert (MoE) (Shazeer et al., 2017), LoRA (Hu et al., 2022), Mixture-of-LoRA (MoL) (Feng et al., 2024), and Task Vector (TV) (Ilharco et al., 2023). MLAR denotes the abbreviation of multi-lingual ability-related corpus.
4.3 The Overall Procedure
To better demonstrate our approach, we present the complete procedure in Algorithm 1 in the pseudo-code form. The procedure of MAET consists of two main stages, i.e., ability-related weights extraction and multi-lingual ability transferring. For the extraction stage, we first utilize the accumulated gradient to estimate the importance of each neuron by Equation 1. Then, we leverage the model trained on the general corpus to remove the influence of language and obtain the ability-specific weight through Equation 2. In the transfer stage, we utilize Equation 3 and Equation 4 to obtain the multi-lingual weight and identify the ability-related parameter tensors in LLM. After it, we leverage Equation 5 to fulfill the multi-lingual ability transfer, to build the multi-lingual ability-enhanced LLM.
To highlight the difference between our approach and previous work, we present the comparison of these methods in Table 1. To adapt LLMs to multi-lingual scenarios, most of the existing methods (e.g., CPT and TV) require the multi-lingual ability-related corpus (i.e., ability-related corpus is required for each language) for training the LLM parameters. In comparison, our proposed approach only trains and modifies the ability-related parameters, which can efficiently focus on enhancing the specific ability. A major novelty of our work is that we identify the key units and implement the sparse update in the model training and merging procedure, which can effectively decompose, extract, and transfer the abilities of LLMs. In addition, compared with the LoRA-based methods (i.e., LoRA and MoL) that also sparsely update the LLM parameters, our approach selectively updates the ability-related neurons, while LoRA-based methods utilize the low-rank matrices to approximate the original parameters.
5 Experiment
5.1 Experimental Settings
In this part, we introduce the details of our experimental settings, including the datasets utilized in the continual pre-training and evaluation process, baseline methods for comparison, and the implementation details of our approach.
Datasets. In this work, we focus on transferring the advanced abilities (i.e., mathematical and scientific reasoning abilities) of LLMs from English scenarios to multi-lingual scenarios, including high-resource languages (i.e., Spanish and Chinese) and low-resource languages (i.e., Bengali and Telugu). For the training corpus, we extract the corpus of the corresponding languages from the dataset proposed by previous work (Yang et al., 2023; Scao et al., 2022; Laurençon et al., 2022) as the general training corpus, and utilize OpenWebMath (Paster et al., 2024) and the arXiv papers (Soldaini et al., 2024) as the ability-related corpus for mathematical tasks and scientific tasks respectively. For the evaluation benchmark, we follow the evaluation settings in previous work (OpenAI, 2023), utilizing Multi-lingual Grade School Math (MGSM) (Shi et al., 2023) and science tasks from multi-lingual MMLU (Lai et al., 2023) (i.e., college biology, college chemistry, college physics, high school biology, high school chemistry, and high school physics) as the downstream tasks for multi-lingual scenarios. The statistical information of the datasets is presented in Table 6.
Baselines. In our evaluation, a baseline can be represented as three parts, i.e., training parameters, training approach, and training data. First, we conduct the full parameters training and the LoRA training (Hu et al., 2022) in our evaluation, denoted as the “F” and “L” at the prefix of the training approaches, respectively. For the training approach, we employ continual pre-training (CPT) (Gururangan et al., 2020), domain adaption (DA) (Taylor et al., 2022), and model merging with task vector (TV) (Ilharco et al., 2023). Besides, for the training data, “L” and “A” refer to the multi-lingual general corpus and the single-lingual ability-related corpus, respectively. Moreover, to conduct a more comprehensive evaluation, we also present the performance of different LLMs, i.e., Baichuan-2 7B (Yang et al., 2023), Mistral 7B (Jiang et al., 2023), LLaMA-2 7B (Touvron et al., 2023), and LLaMA-3 8B (Dubey et al., 2024).
Implementation Details. In the experiment, we adapt LLaMA-3 8B as the backbone LLM, and employ Transformers (Wolf et al., 2020) and Deepspeed  framework to perform the training process. For the training process, the learning rate, batch size, and training step are set as $5\times 10^{-5}$, $1$M tokens, and $2$B tokens, respectively. Besides, for the key neurons locating, we select the top 5% relevant neurons as the key neuron set $\mathcal{N}$ for both stages and identify the last 80% and 60% similar tensor as the key sub-network $\mathcal{T}$ for mathematical reasoning tasks and scientific reasoning tasks respectively. We present model details about hyper-parameters in Appendix B.
5.2 Main Results
To comprehensively evaluate our proposed MAET, we employ MAET on mathematical and scientific tasks in multi-lingual scenarios and present the results in Table 2.
[…]
Last, compared with the model merging based approaches (i.e., F-TV and L-TV), experimental results have shown that MAET performs better than these baseline methods, since we decompose the relation between ability and the language of the training corpus. In the previous model merging approaches, they mainly added the parameters of different models to obtain the final model, without considering the the relation between language and abilities. Due to the extraction mechanism of MAET, we mitigate the effect of languages and make the weight more related to ability, which can be transferred in multi-lingual scenarios.
Methods Multilingual Mathematical Tasks Multilingual Scientific Tasks
ES ZH BN TE Avg. ES ZH BN TE Avg.
Baichuan-2 7B 17.20 28.00 4.80 2.40 13.10 42.27 46.43 30.17 26.21 36.27
Mistral 7B 38.80 34.40 9.60 2.80 21.40 52.08 45.33 32.91 27.96 39.57
LLaMA-2 7B 7.60 12.00 1.60 0.00 5.30 34.16 31.68 24.56 22.15 28.14
LLaMA-3 8B 48.40 38.80 28.80 20.40 34.10 55.06 47.24 36.63 29.26 42.05
[…]
+ F-TV 42.00 32.40 16.00 10.40 25.20 53.36 46.57 36.70 30.73 41.84
+ L-TV 45.60 39.20 30.80 25.60 35.30 55.46 48.27 36.65 30.44 42.71
+ MAET (Ours) 49.60 41.60 32.40 25.20 37.20 56.20 48.00 37.64 30.38 43.06
Table 2: The performance comparison of different approaches on multilingual mathematical tasks and multilingual scientific tasks. Avg. denotes the average accuracy of the multi-lingual tasks. ES, ZH, BN, and TE denote Spanish, Chinese, Bengali, and Telugu, respectively. The best is in bold and the second best is underlined.
Figure 2: The results of ablation study. KNL, AW, API, and AT denote key neurons locating (Eq. 1), ability weights obtaining (Eq. 2), ability-related parameter tensors identifying (Eq. 4), and advanced abilities transferring (Eq. 5).
Figure 3: The performance of different proportions for the ability-related parameters identification.
5.3 Detailed Analysis
To comprehensively evaluate our proposed approach MAET and analyze its features, we conduct several experiments and detailed analysis in this part, including the ablation study, analysis of the transfer ratio of LLM parameters, and the generalization of MAET.
Ablation Study. To assess the effectiveness of each component of our proposed MAET, we conduct the ablation study and present the results in Figure 2. We implement MAET on multi-lingual mathematical and scientific tasks without each module of MAET, i.e., key neurons locating (i.e., Eq. 1), ability weight obtaining (i.e., Eq. 2), ability-related parameter tensor identifying (i.e., Eq. 4), and advanced abilities transferring (Eq. 5). First, in most downstream scenarios, removing any module of MAET will affect the final performance, which has verified the effectiveness of the MAET process. Second, without ability weight obtaining, i.e., directly utilizing the difference between LLM trained on the ability-related corpus and the backbone LLM as the ability weight, we can observe that the performance is seriously hurt in both scenarios, indicating this process can significantly extract the advanced abilities from the single-lingual corpus and decrease the influence of the language of the training corpus. Third, comparing the results of the models whether adopting the ability transferring process, experimental results show that LLM with the multi-lingual abilities enhanced cannot well solve multi-lingual mathematical and scientific tasks, and leveraging the ability weight provided by MAET can improve the LLM performance on advanced tasks.
Ratio of Key Parameters During Transferring Stage. Identifying and updating the ability-related sub-network of LLMs is the key point of our MAET. We conduct experiments to analyze the influence of the transferring ratio $k_{2}\%$ and show the results in Figure 3. Observing the results, the performance of LLM has decreased in a lower and higher ratio of the ability-related parameters identifying process. The main reason is that the lower proportion transfers incomplete knowledge to the model and makes LLM unable to possess the corresponding ability, affecting the performance on the downstream tasks. In contrast, the higher proportion cannot extract the ability weight precisely and will transfer too much language-related knowledge to the model, making the conflict with the LLM inner knowledge and hurting the multi-lingual scenarios performance.
LLM Tensors Proportion of $\mathcal{T}$ ES ZH BN TE Avg.
All Tensors 100.00% 49.60 41.60 32.40 25.20 37.20
Attention All 45.26% 48.80 41.60 28.80 26.40 36.40
   Attention Q 12.07% 47.60 40.80 30.80 26.40 36.40
   Attention K 10.34% 47.20 42.40 29.60 24.40 35.90
   Attention V 9.48% 47.60 42.40 28.80 25.20 36.00
   Attention O 13.36% 48.00 40.40 30.80 27.20 36.60
MLP All 41.38% 48.80 39.60 31.60 27.60 36.90
   MLP Up 13.79% 50.00 40.00 28.80 25.20 36.00
   MLP Gate 13.79% 46.00 41.20 30.00 24.00 35.30
   MLP Down 13.79% 49.60 41.60 30.40 26.00 36.90
Table 3: The effect of only merging the specific LLM tensors during the transferring process (i.e., Eq.5) on multi-lingual mathematical tasks.
Methods MMLU HumanEval MBPP OpenbookQA
LLaMA-3 8B 60.85 35.98 46.60 65.00
+ CPT 58.46 (-2.39) 28.66 (-7.32) 39.60 (-7.00) 61.40 (-3.60)
+ MAET (Ours) 61.07 (+0.22) 35.98 (+0.00) 47.40 (+0.80) 65.00 (+0.00)
Table 4: The out-of-domain performance comparison of different training methods to train LLaMA-3 8B on OpenWebMath. During the ability-enhancing process, previous methods will hurt the OOD abilities of LLM, while our MAET can maintain the corresponding abilities.
Ability-related Sub-networks of LLM. To assess and probe the ability-related sub-networks of LLMs, we only transfer the specific tensors (i.e., tensors in self-attention and MLP mechanism) from the ability weight to the final models through Eq. 5, to analyze the LLM inner abilities. The experimental results are presented in Table 3. From the experiment, we can observe that although the proportion of MLP layers (41.38%) is lower than the attention layers (45.26%), only transferring the MLP layers outperforms transferring the attention layers, indicating that the MLP layers are more related to the advanced abilities and stores the corresponding knowledge. In the MLP layers of LLM, the gate mechanism (i.e., MLP Gate) will control the transmission of information and the down project mechanism (i.e., MLP Down) will integrate the knowledge from previous layers, so that transferring the MLP layers can achieve better performance on the downstream tasks.
Out-of-Domain performance of MAET. We conduct experiments about adapting mathematical ability on the general LLM through MAET, and assess the performance on out-of-domain (OOD) tasks (i.e., MMLU (Hendrycks et al., 2021), HumanEval (Chen et al., 2021a), MBPP (Austin et al., 2021), and OpenbookQA (Mihaylov et al., 2018)), which can reflect and assess different abilities of LLMs. Results are presented in Table 4. We can observe that the performance of LLM on all evaluation tasks has decreased through the CPT training process, and the maximum decrease has been achieved 7.32% on the HumanEval task. One of the possible reasons is that LLaMA-3 has been trained on OpenWebMath during pre-training and the CPT process makes it overfit and forget the knowledge of other domains, hurting the performance on OOD tasks. In contrast, our proposed MAET achieves comparable and even better performance with backbone LLM in all downstream scenarios. Since we identify and update the key neurons related to the specific ability, the ability of LLM can be precisely enhanced, and this strategy also helps the OOD tasks needed for mathematical ability, e.g., mathematical sub-tasks in MMLU and code synthesis task MBPP.
6 Conclusion
In this paper, we presented MAET, which extracted the advanced ability-related weights from the LLM and supported simple addition and subtraction operations to transfer the ability across different languages. Concretely, MAET included two main stages, i.e., extraction and transfer. For the extraction stage, we located the key neurons and extracted the ability-related weights. Then, in the transfer stage, we identified the key parameter tensors and leveraged them to transfer the advanced ability into other LLMs. In this process, the multi-lingual ability-related training corpus is not required, and the experimental results have shown that our approach outperformed competitive baselines.
As future work, we will consider better methods to identify the ability-related sub-network to decompose the abilities of LLMs and utilize an automated approach to determine the hyper-parameter. Besides, we will implement MAET on larger-scale models, and scenarios with more languages and requiring more abilities to evaluate its effectiveness.
7 Limitations
In this section, we discuss the limitations of our work. First, we only implement our approach MAET on 8B LLMs (i.e., LLaMA-3 8B), and do not adopt the LLMs with larger scales (e.g., 13B or 70B LLMs) in the experiment, due to the limitation of computational resources. We will test the effectiveness of our approach on these LLMs in the future. Second, we only evaluate our approach on two downstream tasks (i.e., mathematical and scientific reasoning tasks) in multi-lingual scenarios. Although they are challenging and widely-used testbeds, it is still meaningful to verify our methods on other tasks. Whereas, as we test the performance on diverse high-resource and low-resource languages, it can also provide comprehensive performance estimation for our approach in multi-lingual scenarios. Third, our approach is a general method for ability transferring across different domains, but in this work, we only consider the multi-lingual scenarios and obtain a multi-lingual LLM with the specific ability being enhanced. Forth, our approach also relies on continual pre-training the LLM. Although the training corpus is not very large, it also increases the cost. Fortunately, after we obtain the pre-trained weights, our following steps only need simple addition and subtraction operations for ability transferring, which is flexible for online deployment and application. In future work, we will focus on reducing the data requirement for the pre-training corpus and also improving the weights extracting efficiency. Finally, we do not consider the potential risk and ethics issues that might hurt the alignment of LLMs when using our approach. Actually, our approach is also applicable to transfer the alignment ability across languages. We will investigate to it in the future.
Appendix A Empirical Study
Figure 4: The results of empirical experiments. We present the loss of different training methods during the training process, the cosine similarity of LLM layers after being trained on Zhihu and Reddit, and the similarity of LLMs being trained on different training corpus.
A surge of work (Zhang et al., 2024b; Xiao et al., 2024; Tang et al., 2024) has pointed out that LLMs sparsely activate the specific sub-modules to perform corresponding tasks. Based on these findings, we conduct empirical experiments to explore whether the specific sub-module, which is related to advanced abilities, can be extracted and transferred. We utilize the forum corpus (i.e., Zhihu for Chinese forum corpus and Reddit for English forum corpus) to continually pre-train LLMs, and then assess the training performance (i.e., the value of loss function) and similarity of LLM neurons.
The forum corpus can be considered as containing the question-answering (QA) ability, which is necessary and important for LLMs. The results from Figure 4(a) have shown that only training the top 5% relevant neurons of LLMs can achieve the lower training loss and fit into the training set more quickly, indicating that LLMs contain the sub-module corresponding to the QA ability. Moreover, from Figure 4(b) and Figure 4(c), we can observe that the LLM trained on Zhihu has shown higher similarity with the LLM trained on Reddit than the LLM trained on Github (i.e., lower L1 Norm and higher cosine similarity), and the cosine similarity of different layers in LLM are largely different.
According to the above results, we have found that the different sub-networks of LLMs control the different abilities, and precisely selecting the correct sub-module of LLMs will help the extraction of advanced abilities from the single-lingual corpus and the transfer of these abilities to multi-lingual scenarios. Concretely, although Zhihu and Reddit are in different languages, they will influence the similar sub-modules of LLM and make these sub-networks show high similarity with each other. These sub-networks can be referred to the ability-related sub-networks, which are slightly influenced by languages.
Appendix B Details of Hyper-parameters
We release all of the hyper-parameters during our experiment to better reproduce our proposed approach. Table 5 shows the details of hyper-parameters of different stages.
Stage Hyper-Parameter Mathematical Tasks Scientific Tasks
Extraction Learning Rate $5\times 10^{-5}$ $5\times 10^{-5}$
Batch Size 1M Tokens 1M Tokens
Training Steps 2B Tokens 2B Tokens
$\alpha$ in Extraction 0.8 0.8
$\beta$ in Extraction 0.2 0.2
Ratio of Key Neurons 5% 5%
Extraction Learning Rate $5\times 10^{-5}$ $5\times 10^{-5}$
Batch Size 1M Tokens 1M Tokens
Training Steps 2B Tokens 2B Tokens
$\gamma$ in Transferring 0.2 0.2
$\eta$ in Transferring 1.0 1.0
Ratio of Key Neurons 80% 60%
Table 5: The details of hyper-parameters in the evaluation.
Appendix C Details of Dataset
We present the statistical information of the datasets in Table 6. We mainly consider English, Spanish, Chinese, Bengali, and Telugu in our experiment, and utilized English as the in-domain language while others as the out-of-domain languages. For the evaluation datasets, we select MGSM and multi-lingual MMLU as the evaluation benchmarks, which contain the parallel data in different languages and are useful for multi-lingual complex tasks evaluation.
Language Training Dataset (Tokens) Evaluation Dataset (Instances)
General Corpus Ability-related Corpus Mathematical Tasks Scientific Tasks
English 1.81B 1.30B (Math) / 1.82B (Sci) 250 1,245
Spanish 1.81B - 250 1,232
Chinese 1.80B - 250 1,229
Bengali 1.81B - 250 1,137
Telugu 1.81B - 250 1,036
Table 6: The statistical information of the training and evaluation datasets.

### v2

@@
 In the transfer stage, we further select the ability-related parameter tensors, and design the merging strategy based on the linguistic and ability specific weights, to build the multi-lingual ability-enhanced LLM.
-To demonstrate the effectiveness of our proposed approach, we conduct extensive experiments on mathematical and scientific tasks in both high-resource lingual and low-resource lingual scenarios.
-Experiment results have shown that MAET can effectively and efficiently extract and transfer the advanced abilities, and outperform training-based baselines methods.
+To demonstrate the effectiveness of our proposed approach, we conduct extensive experiments on mathematical and scientific tasks in both high-resource and low-resource lingual scenarios.
+Experiment results have shown that MAET can effectively and efficiently extract and transfer the advanced abilities, outperforming competitive baseline methods and only expending 60% computational resources.
 Our code and data are available at https://github.com/RUCAIBox/MAET.
 1 Introduction
-Large language models (LLMs) have recently shown remarkable performance on various general tasks, e.g., text generation and question answering (Zhao et al., 2023; OpenAI, 2023; Dubey et al., 2024).
-Despite the success, LLMs are still struggling to solve complex tasks (e.g., mathematical reasoning), which require LLMs to possess specific advanced abilities (e.g., deductive reasoning) and knowledge (e.g., mathematical theory) (Yue et al., 2024; Lu et al., 2022).
-To address it and further improve LLMs, existing work either collects the related data to train LLMs (Du et al., 2024; Chen et al., 2024a), or merges the parameters of existing well-performed LLMs to transfer their advanced abilities into one single model (Ilharco et al., 2023; Yadav et al., 2023; Yu et al., 2024a).
+Large language models (LLMs) have recently shown remarkable performance on various general tasks, e.g., text generation and question answering Zhao et al.
+(2023); OpenAI (2023); Dubey et al.
+(2024).
+Despite the success, LLMs are still struggling to solve complex tasks (e.g., mathematical reasoning), which require LLMs to possess specific advanced abilities (e.g., deductive reasoning) and knowledge (e.g., mathematical theory) Yue et al.
+(2024); Lu et al.
+(2022).
+To address it and further improve LLMs, existing work either collects the related data to train LLMs Du et al.
+(2024); Chen et al.
+(2024a), or merges the parameters of existing well-performed LLMs to transfer their advanced abilities into one single model Ilharco et al.
+(2023); Yadav et al.
+(2023); Yu et al.
+(2024a).
 Despite the success, it is not easy to collect sufficient training corpus or well-trained LLMs related to specific abilities, especially in multi-lingual scenarios.
-Especially, some popular languages (e.g., English) have dominated the linguistic expressions of the open web data, and the amount of available data for low-resource languages (e.g., Bengali or Telugu) is highly limited (Patzelt, 2024; Mirashi et al., 2024).
-Fortunately, existing work (Zhao et al., 2024; Schäfer et al., 2024) has revealed that the learned knowledge from one language by LLMs could be inherited and leveraged by other languages.
+Especially, some popular languages (e.g., English) have dominated the linguistic expressions of the open web data, and the amount of available domain-specific data for low-resource languages (e.g., Bengali or Telugu) is highly limited Magueresse et al.
+(2020); Patzelt (2024); Mirashi et al.
+(2024).
+Fortunately, existing work Zhao et al.
+(2024); Schäfer et al.
+(2024) has revealed that the learned knowledge from one language by LLMs could be inherited and leveraged by other languages.
 For example, Llama-series LLMs are trained mainly on English texts, while they can also solve the tasks based on other languages.
-Such a finding has been widely explored in either improving the overall performance of multi-lingual LLMs (Schäfer et al., 2024) or enhancing fine-grained knowledge (Chen et al., 2024a).
+Such a finding has been widely explored in either improving the overall performance of multi-lingual LLMs Schäfer et al.
+(2024) or enhancing fine-grained knowledge Chen et al.
+(2024a).
 However, the related work mostly relies on training with ability-related corpus in the target language, which is not always available for low-resource languages.
@@
 In our approach, we only need ability-specific corpus from any rich-resource language and general multi-lingual corpus, which can effectively mitigate the data scarcity issues in low-resource languages.
-To assess the effectiveness of our approach, we conduct the evaluation based on two comprehensive reasoning benchmarks, namely Multi-lingual Grade School Math (MGSM) (Shi et al., 2023) and science tasks from multi-lingual MMLU (Lai et al., 2023) as the evaluation benchmarks.
-According to the evaluation results, the proposed approach MAET outperforms other competitive baseline methods (e.g., continual pre-training (Gururangan et al., 2020) and model merging methods with task vectors (Ilharco et al., 2023), achieving the 9.1% relative improvement upon the backbone LLM.
+To assess the effectiveness of our approach, we conduct the evaluation based on two comprehensive reasoning benchmarks, namely Multi-lingual Grade School Math (MGSM) Shi et al.
+(2023) and science tasks from multi-lingual MMLU Lai et al.
+(2023) as the evaluation benchmarks.
+According to the evaluation results, the proposed approach MAET outperforms other competitive baseline methods (e.g., continual pre-training Gururangan et al.
+(2020) and model merging methods with task vectors Ilharco et al.
+(2023), achieving the 10% relative improvement compared to the base LLM, which outperforms the competitive baseline methods.
 In addition, our approach can work well with relatively fewer training data, demonstrating an improved efficiency in practice.
-In conclusion, our contribution can be summarized as follows,
-(1) Our research has found that advanced abilities can be extracted from the single-lingual corpus and transferred across languages without the multi-lingual ability-related corpus, enabling to efficiently empower LLMs with special advanced abilities.
-(2) We propose an effective and efficient approach named MAET, which first identifies and extracts the ability-related weights in LLMs and then only leverages simple addition and subtraction operations to build a multi-lingual ability-enhanced LLM.
-(3) Our approach MAET achieves better performance than the competitive baseline methods (e.g., continual pre-training and model merging with task vector) in multi-lingual complex reasoning tasks, including mathematical reasoning tasks and scientific reasoning tasks.
 2 Related Work
-We introduce the related work from the following three perspectives:
 Continual Pre-training.
-Although LLMs have shown remarkable performance on various downstream work, they still struggle in several specific tasks, e.g., complex reasoning tasks (Paster et al., 2024; Shao et al., 2024) or low-resource lingual scenarios (Hedderich et al., 2021; Panchbhai & Pankanti, 2021).
-To adapt LLMs pre-trained on the general corpus to multi-lingual scenarios or specific tasks, existing work (Luo et al., 2022; Taylor et al., 2022; Zhao et al., 2022; Zhang et al., 2024a) has collected the corresponding corpus to continually pre-train (CPT) LLMs.
-During the continual pre-training process, the mixture strategy between the general corpus and the ability-related corpus should be carefully considered to prevent hurting the general abilities of LLMs (Ye et al., 2024; Xie et al., 2023; Chen et al., 2024a; Siriwardhana et al., 2024).
+Although LLMs still struggle in several specific tasks, e.g., complex reasoning tasks Paster et al.
+(2024); Shao et al.
+(2024) or low-resource lingual scenarios Hedderich et al.
+(2021); Panchbhai and Pankanti (2021).
+To adapt LLMs pre-trained on the general corpus to multi-lingual scenarios or specific tasks, existing work Luo et al.
+(2022); Taylor et al.
+(2022); Zhao et al.
+(2022); Zhang et al.
+(2024a) has collected the corresponding corpus to continually pre-train (CPT) LLMs.
+During the continual pre-training process, the mixture strategy between the general corpus and the ability-related corpus should be carefully considered to prevent hurting the general abilities of LLMs Ye et al.
+(2024); Xie et al.
+(2023); Chen et al.
+(2024a); Siriwardhana et al.
+(2024).
 However, previous study Chang et al.
@@
 (2023) has pointed out that it is difficult to collect the required corpus, especially for low-resource language corpus.
-Therefore, synthesizing data from powerful LLMs is widely utilized to expand the task-specific training corpus (Chen et al., 2021b; Yu et al., 2024b; Zhou et al., 2024a).
-Besides, because of the limitation of computation resources, a series of approaches (Hu et al., 2022; Li & Liang, 2021; Dettmers et al., 2023) only train several parameters to reduce the expenses.
+Therefore, synthesizing data from powerful LLMs is widely utilized to expand the task-specific training corpus Chen et al.
+(2021b); Yu et al.
+(2024b); Zhou et al.
+(2024a).
+Besides, because of the limitation of computation resources, a series of approaches Hu et al.
+(2022); Li and Liang (2021); Dettmers et al.
+(2023) only train several parameters to reduce the expenses.
 In this work, we focus on adapting LLMs to multilingual complex reasoning scenarios through continually pre-training LLMs on the single-lingual task-specific corpus.
 Knowledge Editing.
-According to lottery ticket hypothesis (Frankle & Carbin, 2019), training a small number of model parameters will achieve comparable or even better performance on downstream tasks.
-Existing study (Du et al., 2024; Wang et al., 2024b; Gong et al., 2024) has leveraged the inner information of LLMs, e.g., gradient or cosine similarity between different hidden states, to select and train the related sub-network.
-Besides, the probe (i.e., a newly initialized parameter) can be implemented to detect the knowledge of LLMs and process targeted repair (Wang et al., 2024a; Jiang et al., 2024).
-However, Since these approaches need to calculate and select a sub-network of each training instance, which might cause the instability of the training process, several study (Chen et al., 2024b; Zhang et al., 2024b) pointed out that the task-related sub-network can be determined before the training process, and only updating the value of the corresponding neurons can achieve better performance.
-In this work, we focus on editing the task-specific neurons of LLMs to improve the corresponding capacities in multi-lingual scenarios.
+According to lottery ticket hypothesis Frankle and Carbin (2019), training a small number of model parameters will achieve comparable or even better performance on downstream tasks.
+Existing study Du et al.
+(2024); Wang et al.
+(2024b); Gong et al.
+(2024) has leveraged the inner information of LLMs, to select and train the related sub-network.
+Besides, the probe (i.e., a newly initialized parameter) can be implemented to detect the knowledge of LLMs and process targeted repair Wang et al.
+(2024a); Jiang et al.
+(2024).
+However, Since calculating and selecting a sub-network of each training instance might cause the instability of the training process, several study Chen et al.
+(2024b); Zhang et al.
+(2024b) pointed out that the task-related sub-networks can be determined before the training process, and only optimizing these sub-networks can achieve better performance.
+In this work, we focus on editing the task-specific neurons of LLMs to improve the corresponding capacities in multi-lingual tasks.
 Model Merging.
-Given that the CPT process will bring huge computational expenses, previous work leveraged model merging techniques to integrate different abilities (e.g., mathematical reasoning and code synthesizing) into one model (Yang et al., 2024; Xu et al., 2024; Stoica et al., 2024).
+Given that the CPT process will bring huge computational expenses, previous work leveraged model merging techniques to integrate different abilities (e.g., mathematical reasoning and code synthesizing) into one model Yang et al.
+(2024); Xu et al.
+(2024b); Stoica et al.
+(2024).
 During the merging process, the interference between different LLMs might be conflict with each other and affect the final performance.
-Therefore, researchers proposed the clip (Yadav et al., 2023) or randomly dropout (Yu et al., 2024a) mechanism to mitigate the performance decrease.
-Moreover, the selection of the hyper-parameters (e.g., weight of each model) is the challenge of the model merging process, and previous work (Zhou et al., 2024b; Matena & Raffel, 2022) utilized the inner parameters of LLMs or external matrixes to determine the hyper-parameters.
+Therefore, researchers proposed the clip Yadav et al.
+(2023) or randomly dropout Yu et al.
+(2024a) mechanism to mitigate the performance decrease.
+Moreover, the selection of the hyper-parameters (e.g., weight of each model) is the challenge of the model merging process, and previous work Zhou et al.
+(2024b); Matena and Raffel (2022) utilized the inner parameters of LLMs or external matrixes to determine the hyper-parameters.
+Furthermore, a series of work has studied improving the reasoning ability of LLMs in non-English scenarios by merging the reasoning-specialized model and multi-lingual model Huang et al.
+(2024); Yoon et al.
+(2024).
 Inspired by the above work, we try to locate the task-related sub-networks of LLMs and transfer the advanced abilities.
 3 Preliminary
-Despite that LLMs exhibit excellent performance on general tasks, they still have limited advanced abilities, e.g., mathematical and scientific reasoning abilities.
-A typical approach to enhancing these abilities is to continually pre-train (CPT) LLMs with ability-related corpus.
+Despite that LLMs exhibit remarkable performance on general tasks, they still have limited advanced abilities, e.g., mathematical and scientific reasoning abilities.
+A typical approach to enhance these abilities is to continually pre-train (CPT) LLMs with ability-related corpus.
 However, such training data might not always be available or sufficient, especially for minor domains (e.g., Bengali).
 In this work, we focus on the task of ability extraction and transfer by continual pre-training and merging LLMs.
-Concretely, LLMs are trained on collected corpus from a certain domain, and we aim to only transfer its learned advanced capabilities to another target domain (Zhuang et al., 2021; Farahani et al., 2021) without further training efforts.
-In this work, we mainly study the cross-lingual scene where the linguistic-agnostic advanced ability and linguistic abilities should be extracted and transferred, to help build a unified multi-lingual ability-enhanced LLM.
+Concretely, LLMs are trained on the collected corpus from a certain domain, and we aim to only transfer its learned advanced capabilities to target domains (Zhuang et al., 2021; Farahani et al., 2021) without further training.
+In this work, we study the cross-lingual scene where the linguistic-agnostic advanced ability and linguistic abilities should be extracted and transferred, to build a unified multi-lingual ability-enhanced LLM.
 Formally, for a certain ability $A_{i}$ and a set of languages $L=\{L_{0},L_{1},\dots,L_{n}\}$, we assume that the general corpus of all languages can be collected, denoted as $\mathcal{C}_{\text{general}}=\{\mathcal{C}_{L_{0}},\mathcal{C}_{L_{1}},\dots,\mathcal{C}_{L_{n}}\}$, while the ability-related corpus is only available in language $L_{0}$ (i.e., English), denoted as $\mathcal{C}_{L_{0},A_{i}}$.
@@
 Figure 1: The framework of our approach MAET, including extraction and transfer stages.
-In the extraction stage, we first locate the key neurons, and utilize the single-lingual ability-related corpus and general corpus to train the LLM on these neurons to obtain the ability-specific weight.
+In the extraction stage, we first locate the key neurons, and utilize the single-lingual ability-related corpus and general corpus to train the LLM on these neurons to obtain the ability-related weight.
 Then, we remove the parameter tensors related to language knowledge in the ability weight and transfer the remaining to the base LLM.
@@
 The key motivation of our approach is to identify and extract ability-related neurons or weights, and transfer the target abilities into a LLM in an efficient way.
-The overall framework of MAET is presented in Figure 1.
+The framework of MAET is presented in Figure 1.
 4.1 Ability-related Weights Extraction
@@
 Locating the Key Neurons.
-The gradient of each neuron in LLMs can be utilized to estimate its correlation degree with specific task ability , we select those with high gradient values as key neurons.
-To this end, we first use a small amount of ability-related corpus $\mathcal{C}_{L_{0},A_{i}}^{\prime}$ to continually pre-train the LLM.
+The gradient of each neuron in LLMs can be utilized to estimate its correlation degree with specific task ability Pruthi et al.
+(2020); Chen et al.
+(2024b); Xia et al.
+(2024), we select those with high gradient values as key neurons.
+To this end, we first use the ability-related corpus $\mathcal{C}_{L_{0},A_{i}}^{\prime}$ to continually pre-train the LLM, while sampling a small amount to train the model can be also applied to reduce the computation consumption.
 During training, the LLM learns the language modeling task and each neuron is updated by the gradients associated by the training instances.
@@
 Formally, the importance function $I(A_{i},\theta_{j})$ of neurons can be computed as:
-$I(A_{i},\theta_{j})=\sum_{d_{k}\in\mathcal{C}_{L_{0},A_{i}}^{\prime}}{\text{Gradient}\left(\theta_{j},d_{k}\right)}\approx\lambda\cdot\parallel\tilde{\theta}_{j}-\theta_{j}\parallel,$ (1)
+$I(A_{i},\theta_{j})=\sum_{d_{k}\in\mathcal{C}_{L_{0},A_{i}}^{\prime}}{\text{Grad}\left(\theta_{j},d_{k}\right)}\approx\frac{\parallel\tilde{\theta}_{j}-\theta_{j}\parallel}{\text{LearningRate}},$ (1)
 where $d_{k}$ denotes the $k$-th instance of training corpus $\mathcal{C}_{L_{0},A_{i}}^{\prime}$ and $\tilde{\theta}_{j}$ denote the value of the $j$-th neuron of LLM after training, respectively.
-Based on it, we rank all neurons according to their importance scores, and then select the top $k_{1}\%$ ones into the set $\mathcal{N}_{A_{i}}$ as the key neurons.
-Learning Ability-specific Weights.
-Based on the identified key neurons in $\mathcal{N}_{A_{i}}$, we further learn the ability-specific parameter weights.
+Based on it and inspired by previous work (Yadav et al., 2023), we rank all neurons according to their importance scores, and then select the top $k_{1}\%$ ones into the set $\mathcal{N}_{A_{i}}$ as the key neurons.
+Learning Ability-related Weights.
+Based on the identified key neurons in $\mathcal{N}_{A_{i}}$, we further learn the ability-related parameter weights.
 Our motivation is to decompose the parameter weights according to their changes before and after the LLM has mastered a specific ability, which is achievable owing to the modularity and composition nature of the LLM parameter matrices (Yu et al., 2024a; Shazeer et al., 2017).
@@
 Then, we train the LLM with the mixture of ability-related corpus and general corpus on the key neuron set $\mathcal{N}_{A_{i}}\bigcup\mathcal{N}_{L_{0}}$ and $\mathcal{N}_{L_{0}}$ respectively, to obtain two specific models, denoted as $\text{LLM}_{A_{i},L_{0}}$ with parameters $\Theta_{A_{i},L_{0}}$ and $\text{LLM}_{L_{0}}$ with parameters $\Theta_{L_{0}}$.
-Next, we measure the parameter changes between the backbone and the trained models, and obtain the ability-specific weights via the parameter decomposition operation as:
+Next, we measure the parameter changes between the backbone and the trained models, and obtain the ability-related weights via the parameter decomposition operation as:
 $R(A_{i})=\alpha\cdot\underbrace{(\Theta_{A_{i},L_{0}}-\Theta_{o})}_{\text{Ability \& language difference}}-\beta\cdot\underbrace{(\Theta_{L_{0}}-\Theta_{o})}_{\text{Language difference}},$ (2)
@@
 As we only train the parameters within the neuron set, its weight difference should preserve the knowledge about the corresponding ability.
-Thus, it can be regarded as the ability-specific parameter representations, and is promising to transfer the ability into other LLMs by the addition operation.
-1
-Input : Single-lingual ability-related corpus $\mathcal{C}_{L_{0},A_{i}}$, multi-lingual general corpus $\mathcal{C}_{L_{0}},\mathcal{C}_{L_{1}},\dots,\mathcal{C}_{L_{n}}$, and the parameters of the backbone model $\Theta_{o}$.
-Output : A well-trained multi-lingual ability-enhanced LLM.
-2
-// Ability-related Weights Extraction
-3 $\theta^{\prime}\leftarrow\text{CPT}(\mathcal{C}_{L_{0},A_{i}},\Theta_{o}$);
-4
-5for $j$-th neuron in $\Theta_{o}$ do
-6       Calculate the importance score of the corresponding neuron using Eq. 1;
-7
-8
-9Identify the key neuron set $\mathcal{N}_{A_{i}}$;
-10
-[…]
-12
-[…]
-14
-15Learning the ability-specific weight $R(A_{i})$ using Eq. 2;
-16
-// Multi-lingual Ability Transfer
-17
-18Obtaining the multi-lingual weight $R_{Lang}$ using Eq. 3;
-19
-20for $j$-th parameter tensor in LLM do
-21       Calculate the correlation using Eq. 4;
-22
-23
-24Identify the ability-related parameters $\mathcal{T}$;
-25
-26Transfer the ability to multi-lingual scenarios using Eq. 5;
-27
-28Obtain the well-trained multi-lingual ability-enhanced LLM.
-Algorithm 1 The complete procedure of our proposed approach.
+Thus, it can be regarded as the ability-related parameter representations, and is promising to transfer the ability into other LLMs by the addition operation.
 4.2 Multi-lingual Ability Transfer
-After obtaining the ability-related weights, in this part, we utilize them to transfer and integrate the abilities, to build a multi-lingual ability-enhanced LLM.
+After obtaining the ability-related weights, we utilize them to transfer and integrate the abilities, to build a multi-lingual ability-enhanced LLM.
 Ability-related Parameter Tensor Selection.
@@
 $S(\tau_{i})=\texttt{sim}\left(R(A_{i})[\tau_{i}],R_{Lang}[\tau_{i}]\right),$ (4)
-where we use the dot product to implement the similarity function $\texttt{sim}\left(\cdot\right)$.
+where we use the cosine similarity to implement the similarity function $\texttt{sim}\left(\cdot\right)$.
 After obtaining the similarity of all tensors, we rank them in a descending order based on the similarity values, and then select the last $k_{2}\%$ parameters into the set $\mathcal{T}$ as the ability-related parameters.
@@
 In this way, we can derive a more powerful LLM that is equipped with the multi-lingual abilities and specific advanced abilities.
-Aspects CPT MoE LoRA MoL TV MAET (Ours)
-MLAR Corpus Yes Yes Yes Yes Yes No
-Tuning Parameters Full Full Low-Rank Low-Rank Full Ability-related
-Ability Composition No No No No Yes Yes
-Ability Transfer No No No No No Yes
-Table 1: The difference between our proposed MAET and the methods in previous work (i.e., CPT (Hu et al., 2022), Mixture-of-Expert (MoE) (Shazeer et al., 2017), LoRA (Hu et al., 2022), Mixture-of-LoRA (MoL) (Feng et al., 2024), and Task Vector (TV) (Ilharco et al., 2023).
-MLAR denotes the abbreviation of multi-lingual ability-related corpus.
+Approaches MLAR TPara AC AT
+CPT Yes Full No No
+MoE Yes Full No No
+LoRA Yes Low-Rank No No
+MoL Yes Low-Rank No No
+TV Yes Full Yes No
+MAET No Ability-related Yes Yes
+Table 1: The difference between our MAET and the methods in previous work (i.e., CPT (Hu et al., 2022), Mixture-of-Expert (MoE) (Shazeer et al., 2017), LoRA (Hu et al., 2022), Mixture-of-LoRA (MoL) (Feng et al., 2024), and Task Vector (TV) (Ilharco et al., 2023).
+MLAR, TPara, AC, and AT denote the abbreviation of multi-lingual ability-related corpus, parameters for training, ability composition, and ability transfer.
 4.3 The Overall Procedure
-To better demonstrate our approach, we present the complete procedure in Algorithm 1 in the pseudo-code form.
+To better demonstrate our approach, we present key concepts in Table 4 for further clarifying and provide the complete procedure in Algorithm 1 in the pseudo-code form.
 The procedure of MAET consists of two main stages, i.e., ability-related weights extraction and multi-lingual ability transferring.
 For the extraction stage, we first utilize the accumulated gradient to estimate the importance of each neuron by Equation 1.
-Then, we leverage the model trained on the general corpus to remove the influence of language and obtain the ability-specific weight through Equation 2.
+Then, we leverage the model trained on the general corpus to remove the influence of language and obtain the ability-related weight through Equation 2.
 In the transfer stage, we utilize Equation 3 and Equation 4 to obtain the multi-lingual weight and identify the ability-related parameter tensors in LLM.
@@
 5.1 Experimental Settings
-In this part, we introduce the details of our experimental settings, including the datasets utilized in the continual pre-training and evaluation process, baseline methods for comparison, and the implementation details of our approach.
+We introduce the datasets and the baseline methods in our evaluation in this part, and present the implementation details of our approach in Appendix B.
 Datasets.
-In this work, we focus on transferring the advanced abilities (i.e., mathematical and scientific reasoning abilities) of LLMs from English scenarios to multi-lingual scenarios, including high-resource languages (i.e., Spanish and Chinese) and low-resource languages (i.e., Bengali and Telugu).
-For the training corpus, we extract the corpus of the corresponding languages from the dataset proposed by previous work (Yang et al., 2023; Scao et al., 2022; Laurençon et al., 2022) as the general training corpus, and utilize OpenWebMath (Paster et al., 2024) and the arXiv papers (Soldaini et al., 2024) as the ability-related corpus for mathematical tasks and scientific tasks respectively.
-For the evaluation benchmark, we follow the evaluation settings in previous work (OpenAI, 2023), utilizing Multi-lingual Grade School Math (MGSM) (Shi et al., 2023) and science tasks from multi-lingual MMLU (Lai et al., 2023) (i.e., college biology, college chemistry, college physics, high school biology, high school chemistry, and high school physics) as the downstream tasks for multi-lingual scenarios.
-The statistical information of the datasets is presented in Table 6.
+We focus on transferring the advanced abilities (i.e., mathematical and scientific reasoning abilities) of LLMs from English scenarios to multi-lingual scenarios, including high-resource languages (i.e., Spanish) and low-resource languages (i.e., Bengali and Telugu).
+Thus, for the training corpus, we extract the corpus proposed by previous work (Yang et al., 2023; Scao et al., 2022; Laurençon et al., 2022) as the general training corpus, and utilize OpenWebMath (Paster et al., 2024) and the arXiv papers (Soldaini et al., 2024) as the ability-related corpus for mathematical tasks and scientific tasks, respectively.
+For the evaluation benchmark, we follow the evaluation settings in previous work (OpenAI, 2023), utilizing Multi-lingual Grade School Math (MGSM) (Shi et al., 2023) and science tasks from multi-lingual MMLU (Lai et al., 2023) (i.e., college biology, college chemistry, college physics, high school biology, high school chemistry, and high school physics) as the downstream tasks.
+The statistical information of the datasets is presented in Table 6 in the Appendix.
 Baselines.
-In our evaluation, a baseline can be represented as three parts, i.e., training parameters, training approach, and training data.
-First, we conduct the full parameters training and the LoRA training (Hu et al., 2022) in our evaluation, denoted as the “F” and “L” at the prefix of the training approaches, respectively.
-For the training approach, we employ continual pre-training (CPT) (Gururangan et al., 2020), domain adaption (DA) (Taylor et al., 2022), and model merging with task vector (TV) (Ilharco et al., 2023).
-Besides, for the training data, “L” and “A” refer to the multi-lingual general corpus and the single-lingual ability-related corpus, respectively.
-Moreover, to conduct a more comprehensive evaluation, we also present the performance of different LLMs, i.e., Baichuan-2 7B (Yang et al., 2023), Mistral 7B (Jiang et al., 2023), LLaMA-2 7B (Touvron et al., 2023), and LLaMA-3 8B (Dubey et al., 2024).
-Implementation Details.
-In the experiment, we adapt LLaMA-3 8B as the backbone LLM, and employ Transformers (Wolf et al., 2020) and Deepspeed  framework to perform the training process.
-For the training process, the learning rate, batch size, and training step are set as $5\times 10^{-5}$, $1$M tokens, and $2$B tokens, respectively.
-Besides, for the key neurons locating, we select the top 5% relevant neurons as the key neuron set $\mathcal{N}$ for both stages and identify the last 80% and 60% similar tensor as the key sub-network $\mathcal{T}$ for mathematical reasoning tasks and scientific reasoning tasks respectively.
-We present model details about hyper-parameters in Appendix B.
-5.2 Main Results
-To comprehensively evaluate our proposed MAET, we employ MAET on mathematical and scientific tasks in multi-lingual scenarios and present the results in Table 2.
-[…]
-[…]
-[…]
-Last, compared with the model merging based approaches (i.e., F-TV and L-TV), experimental results have shown that MAET performs better than these baseline methods, since we decompose the relation between ability and the language of the training corpus.
-In the previous model merging approaches, they mainly added the parameters of different models to obtain the final model, without considering the the relation between language and abilities.
-Due to the extraction mechanism of MAET, we mitigate the effect of languages and make the weight more related to ability, which can be transferred in multi-lingual scenarios.
-Methods Multilingual Mathematical Tasks Multilingual Scientific Tasks
-ES ZH BN TE Avg. ES ZH BN TE Avg.
-Baichuan-2 7B 17.20 28.00 4.80 2.40 13.10 42.27 46.43 30.17 26.21 36.27
-Mistral 7B 38.80 34.40 9.60 2.80 21.40 52.08 45.33 32.91 27.96 39.57
-LLaMA-2 7B 7.60 12.00 1.60 0.00 5.30 34.16 31.68 24.56 22.15 28.14
-LLaMA-3 8B 48.40 38.80 28.80 20.40 34.10 55.06 47.24 36.63 29.26 42.05
+In our evaluation, we utilize LLaMA-3 8B (Dubey et al., 2024) as the backbone model and adopt four categories of the widely used methods as baselines, i.e., continually pre-training, transfer learning, data augmentation, and model merging based approaches.
+Concretely, a baseline can be represented as three parts, i.e., training parameters, training approach, and training data.
+First, we conduct the full parameters training and the LoRA training (Hu et al., 2022), which are denoted as the “F” and “L” at the prefix, respectively.
+Second, for the training approach, we employ continual pre-training (CPT) (Gururangan et al., 2020), domain adaption (DA) (Taylor et al., 2022), and model merging with task vector (TV) (Ilharco et al., 2023).
+Third, for the training data, “L”, “A”, and “T” refer to the multi-lingual general corpus, English ability-related corpus, and multi-lingual ability-related corpus translated from GPT-4o (Hurst et al., 2024), respectively.
+Moreover, we also present the performance of other LLMs, i.e., Baichuan-2 7B (Yang et al., 2023), Mistral 7B (Jiang et al., 2023), and LLaMA-2 7B (Touvron et al., 2023).
+Methods #Tokens Multilingual Mathematical Tasks Multilingual Scientific Tasks
+ES BN TE Avg. RICER ($\uparrow$) ES BN TE Avg. RICER ($\uparrow$)
+Baichuan-2 7B - 17.2 4.8 2.4 8.1 - 42.3 30.2 26.2 32.9 -
+Mistral 7B - 38.8 9.6 2.8 17.1 - 52.1 32.9 28.0 37.7 -
+LLaMA-2 7B - 7.6 1.6 0.0 3.1 - 34.2 24.6 22.2 27.0 -
+LLaMA-3 8B - 48.4 28.8 20.4 32.5 - 55.1 36.6 29.3 40.3 -
+Vanilla Continually Pre-training based Approaches
 […]
@@
 […]
+Transfer Learning based Approaches
 […]
-+ F-TV 42.00 32.40 16.00 10.40 25.20 53.36 46.57 36.70 30.73 41.84
-+ L-TV 45.60 39.20 30.80 25.60 35.30 55.46 48.27 36.65 30.44 42.71
-+ MAET (Ours) 49.60 41.60 32.40 25.20 37.20 56.20 48.00 37.64 30.38 43.06
-Table 2: The performance comparison of different approaches on multilingual mathematical tasks and multilingual scientific tasks.
-Avg.
-denotes the average accuracy of the multi-lingual tasks.
-ES, ZH, BN, and TE denote Spanish, Chinese, Bengali, and Telugu, respectively.
-The best is in bold and the second best is underlined.
-Figure 2: The results of ablation study.
-KNL, AW, API, and AT denote key neurons locating (Eq.
+[…]
+Data Augmentation based Approaches
+[…]
+[…]
+Model Merging based Approaches
++ F-TV 12B 42.0 16.0 10.4 22.8 -80.8 53.4 36.7 30.7 40.3 -0.3
++ L-TV 12B 45.6 30.8 25.6 34.0 12.5 55.5 36.7 30.4 40.9 4.6
++ MAET (Ours) 12B 49.6 32.4 25.2 35.7 26.9 56.2 37.6 30.4 41.4 9.2
+Table 2: The performance of different approaches on multilingual mathematical and scientific tasks.
+ES, BN, and TE denote Spanish, Bengali, and Telugu, respectively.
+#Tokens denotes the number of training tokens.
+RICER refers to the reciprocal of incremental cost effectiveness ratio, i.e., RICER = #Tokens / Improvement ×$\times 100\%$.
+5.2 Main Results
+We present the evaluation results of MAET and the baseline methods in Table 2.
+[…]
+Second, although our MAET shows similar training efficiency to transfer learning based approaches, MAET performs better than these baseline methods, showing the higher RICER score (e.g., 26.9 v.s.
+16.9).
+For transfer learning based approaches, since the model is only trained on the single-lingual ability-related corpus during the domain adaptation process, it is difficult for LLM to handle the challenging tasks in multi-lingual scenarios.
+Concretely, the performance of LLaMA-3 8B on the multi-lingual scientific tasks even decreases after domain adaptation, showing a 4% relative decrease.
+To alleviate this issue, MAET leverages the calculation between the parameters of different models to extract the ability-related weights, which are language-agnostic and can be transferred to any other scenario.
+Third, MAET also achieves higher performance than data augmentation based approaches (i.e., training LLM on the multi-lingual ability-related corpus translated by GPT-4o).
+The translation-based method consumes more computational resources and cannot achieve better performance.
+The reason might be that LLMs cannot perform the translation process well and the translated corpus shares similar knowledge of the specific domain, which makes LLM overfit the corresponding knowledge and cannot really understand the specific knowledge.
+In contrast, our approach decomposes the advanced ability and language ability, and transfers the advanced ability from one language to another, preventing overfitting, decreasing the expense, and improving performance.
+These results demonstrate that data-centric methods are difficult to build a multi-lingual ability-enhanced LLM.
+Last, compared with the model merging based approaches (i.e., F-TV and L-TV), experimental results have shown that MAET performs better than these baseline methods, since we decompose the relation between ability and the language of the training corpus.
+In the previous model merging approaches, they mainly added the parameters of different models to obtain the final model, without considering the the relation between language and abilities.
+Due to the extraction mechanism of MAET, we mitigate the effect of languages and make the weight more related to ability, which can be transferred in multi-lingual scenarios.
+5.3 Detailed Analysis
+To further analyze the features of MAET, we conduct the detailed analysis in this part, including the ablation study, analysis of the transfer ratio $k_{2}$, and the generalization of MAET.
+Figure 2: The ablation study.
+KNL, AWO, ATI, and AAT denote key neurons locating (Eq.
 1), ability weights obtaining (Eq.
-2), ability-related parameter tensors identifying (Eq.
+2), ability-related tensors identifying (Eq.
 4), and advanced abilities transferring (Eq.
 5).
-Figure 3: The performance of different proportions for the ability-related parameters identification.
-5.3 Detailed Analysis
-To comprehensively evaluate our proposed approach MAET and analyze its features, we conduct several experiments and detailed analysis in this part, including the ablation study, analysis of the transfer ratio of LLM parameters, and the generalization of MAET.
 Ablation Study.
@@
 Identifying and updating the ability-related sub-network of LLMs is the key point of our MAET.
-We conduct experiments to analyze the influence of the transferring ratio $k_{2}\%$ and show the results in Figure 3.
+We analyze the influence of the transferring ratio $k_{2}\%$ and show the results in Figure 3.
 Observing the results, the performance of LLM has decreased in a lower and higher ratio of the ability-related parameters identifying process.
 The main reason is that the lower proportion transfers incomplete knowledge to the model and makes LLM unable to possess the corresponding ability, affecting the performance on the downstream tasks.
-In contrast, the higher proportion cannot extract the ability weight precisely and will transfer too much language-related knowledge to the model, making the conflict with the LLM inner knowledge and hurting the multi-lingual scenarios performance.
+In contrast, the higher proportion cannot extract the ability weight precisely and will transfer too much language-related knowledge to the model, making the conflict with the LLM’s inner knowledge and hurting the multi-lingual scenarios’ performance.
+Figure 3: The performance of different proportions for the ability-related parameters identification.
+Out-of-Domain Performance of MAET.
+We conduct experiments about adapting mathematical ability on LLaMA-3 8B through MAET, and assess its performance on out-of-domain (OOD) tasks (i.e., MMLU (Hendrycks et al., 2021), HumanEval (Chen et al., 2021a), MBPP (Austin et al., 2021), and OpenbookQA (Mihaylov et al., 2018)).
+Results are presented in Table 3.
+We can observe that the performance of LLM on all evaluation tasks has decreased through the CPT training process, and the maximum decrease has been achieved 7.32% on the HumanEval task.
+One of the possible reasons is that LLaMA-3 has been trained on OpenWebMath during pre-training and the CPT process makes it overfit and forget the knowledge of other domains, hurting the performance on OOD tasks.
+In contrast, our proposed MAET achieves comparable and even better performance with backbone LLM in all downstream scenarios.
+Since we identify and update the key neurons related to the specific ability, the ability of LLM can be precisely enhanced, and this strategy also helps the OOD tasks needed for mathematical ability, e.g., mathematical tasks in MMLU and MBPP.
+Methods MMLU MBPP OpenbookQA
+LLaMA-3 8B 60.85 46.60 65.00
++ CPT -2.39 -7.00 -3.60
++ MAET +0.22 +0.80 +0.00
+Table 3: The out-of-domain performance of different training methods to train LLaMA-3 8B on OpenWebMath.
+During the ability-enhancing process, previous methods will hurt the OOD abilities of LLM, while our MAET can maintain the corresponding abilities.
+6 Conclusion
+In this paper, we presented MAET, which extracted the advanced ability-related weights from the LLM and supported simple addition and subtraction operations to transfer the ability across different languages.
+Concretely, MAET included two main stages, i.e., extraction and transfer.
+For the extraction stage, we located the key neurons and extracted the ability-related weights.
+Then, in the transfer stage, we identified the key parameter tensors and leveraged them to transfer the advanced ability into other LLMs.
+In this process, the multi-lingual ability-related training corpus is not required, and the experimental results have shown that our approach outperformed competitive baselines.
+As future work, we will consider better methods to identify the ability-related sub-network to decompose the abilities of LLMs and utilize an automated approach to determine the hyper-parameter.
+Besides, we will implement MAET on larger-scale models, and scenarios with more languages and requiring more abilities to evaluate its effectiveness.
+Limitations
+In this section, we discuss the limitations of our work.
+First, we only implement our approach MAET on 8B LLMs (i.e., LLaMA-3 8B), and do not adopt the LLMs with larger scales (e.g., 13B or 70B LLMs) in the experiment, due to the limitation of computational resources.
+We will test the effectiveness of our approach on these LLMs in the future.
+Second, we only evaluate our approach on two downstream tasks (i.e., mathematical and scientific reasoning tasks) in multi-lingual scenarios.
+Although they are challenging and widely-used testbeds, it is still meaningful to verify our methods on other tasks.
+Whereas, as we test the performance on diverse high-resource and low-resource languages, it can also provide comprehensive performance estimation for our approach in multi-lingual scenarios.
+Finally, we do not consider the potential risk and ethics issues that might hurt the alignment of LLMs when using our approach.
+Actually, our approach is also applicable to transfer the alignment ability across languages.
+We will investigate to it in the future.
+Appendix A Empirical Study
+A surge of work (Zhang et al., 2024b; Xiao et al., 2024; Tang et al., 2024) has pointed out that LLMs sparsely activate the specific sub-modules to perform corresponding tasks.
+Based on these findings, we conduct empirical experiments to explore whether the specific sub-module, which is related to advanced abilities, can be extracted and transferred.
+We utilize the forum corpus (i.e., Zhihu for Chinese forum corpus and Reddit for English forum corpus) to continually pre-train LLMs, and then assess the training performance (i.e., the value of loss function) and similarity of LLM neurons.
+The forum corpus can be considered as containing the question-answering (QA) ability, which is necessary and important for LLMs.
+The results from Figure 4(a) have shown that only training the top 5% relevant neurons of LLMs can achieve the lower training loss and fit into the training set more quickly, indicating that LLMs contain the sub-module corresponding to the QA ability.
+Moreover, from Figure 4(b) and Figure 4(c), we can observe that the LLM trained on Zhihu has shown higher similarity with the LLM trained on Reddit than the LLM trained on Github (i.e., lower L1 Norm and higher cosine similarity), and the cosine similarity of different layers in LLM are largely different.
+According to the above results, we have found that the different sub-networks of LLMs control the different abilities, and precisely selecting the correct sub-module of LLMs will help the extraction of advanced abilities from the single-lingual corpus and the transfer of these abilities to multi-lingual scenarios.
+Concretely, although Zhihu and Reddit are in different languages, they will influence the similar sub-modules of LLM and make these sub-networks show high similarity with each other.
+These sub-networks can be referred to the ability-related sub-networks, which are slightly influenced by languages.
+Stage Hyper-Parameter Mathematical Tasks Scientific Tasks
+Extracting Learning Rate $5\times 10^{-5}$ $5\times 10^{-5}$
+Batch Size 1M Tokens 1M Tokens
+Training Steps 2B Tokens 2B Tokens
+$\alpha$ in Extraction 0.8 0.8
+$\beta$ in Extraction 0.2 0.2
+Ratio of Key Neurons 5% 5%
+Transferring Learning Rate $5\times 10^{-5}$ $5\times 10^{-5}$
+Batch Size 1M Tokens 1M Tokens
+Training Steps 2B Tokens 2B Tokens
+$\gamma$ in Transferring 0.2 0.2
+$\eta$ in Transferring 1.0 1.0
+Ratio of Key Neurons 80% 60%
+$\mu$ for Spanish 1.5 1.5
+$\mu$ for Chinese 2.0 2.0
+$\mu$ for Bengali 1.2 1.2
+$\mu$ for Telugu 1.2 1.2
+Table 5: The details of hyper-parameters in the training and evaluation process.
+Language Training Dataset (Tokens) Evaluation Dataset (Instances)
+General Corpus Ability-related Corpus Mathematical Tasks Scientific Tasks
+English 1.81B 1.30B (Math) / 1.82B (Sci) 250 1,245
+Spanish 1.81B - 250 1,232
+Chinese 1.80B - 250 1,229
+Bengali 1.81B - 250 1,137
+Telugu 1.81B - 250 1,036
+Table 6: The statistical information of the training and evaluation datasets.
+Methods Qwen2.5 0.5B Gemma2 2B
+ES TE Avg. ES TE Avg.
+Backbone LLM 36.64 25.69 31.17 43.41 30.01 36.71
+[…]
+[…]
++ MAET w/o API 36.72 28.91 32.82 43.23 29.59 36.41
++ MAET (Ours) 36.91 29.62 33.27 43.62 30.37 37.00
+Table 7: The performance comparison of different LLMs on multilingual scientific tasks.
+Appendix B Implementation Details
+In the experiment, we adapt LLaMA-3 8B as the backbone LLM, and employ Transformers (Wolf et al., 2020) and Deepspeed  framework to perform the training process.
+And we also present the evaluation results of different backbone LLM (i.e., Qwen2.5 0.5B (Hui et al., 2024) and Gemma2 2B (Rivière et al., 2024)) in Appendix E.
+For the training process, the learning rate, batch size, and training step are set as $5\times 10^{-5}$, $1$M tokens, and $2$B tokens, respectively.
+Besides, for the key neurons locating, we select the top 5% relevant neurons as the key neuron set $\mathcal{N}$ for both stages and identify the last 80% and 60% similar tensor as the key sub-network $\mathcal{T}$ for mathematical reasoning tasks and scientific reasoning tasks respectively.
+Hyper-parameters Selection.
+we released all of the hyper-parameters during our experiment in Table 5, to reproduce our proposed approach better.
+The hyperparameters discussed in the paper can be categorized into two types: training-related parameters (e.g., learning rate, batch size) and training-independent parameters (i.e., $\alpha$, $\beta$, $\gamma$, $\eta$, and $\mu$).
+Training-related parameters do not require extensive hyperparameter tuning, as existing studies Dubey et al.
+(2024); Hui et al.
+(2024) provide clear guidelines for setting them.
+On the other hand, training-independent parameters are used to construct ability-related weights, tensors, and language-specific weights.
+These techniques are similar to those employed in model merging Ilharco et al.
+(2023); Yadav et al.
+(2023), and the hyperparameter setting approach outlined in the paper can be applied.
+A limited number of hyperparameter sets can be defined and validated, as the process primarily involves simple additions and subtractions of model parameters, making it computationally inexpensive.
+Appendix C Details of Dataset
+We present the statistical information of the datasets in Table 6.
+We mainly consider English, Spanish, Chinese, Bengali, and Telugu in our experiment, and utilized English as the in-domain language while others as the out-of-domain languages.
+For the evaluation datasets, we select MGSM and multi-lingual MMLU as the evaluation benchmarks, which contain the parallel data in different languages and are useful for multi-lingual complex tasks evaluation.
+Appendix D Prompt for Translation
+You should translate the following text from English to {TARGET LANGUAGE} and should not modify the latex code or website code.
+You should not add any details that are not mentioned in the original text.
+## English
+{ENGLISH TEXT}
+## {TARGET LANGUAGE}
+Appendix E Performance of Small Scale LLMs
+We conduct the different LLMs with different sizes (i.e.,, Qwen2.5-0.5B and Gemma2-2B) in our experiment to valid the practicality of our approach.
+We assess MAET and baselines on multi-lingual scientific reasoning tasks and present the evaluation results in Table 7.
+Comparing the performance of MAET and the baseline methods, we can observe that MAET can also enhance the performance of small scale models and outperform competitive baselines.
+Therefore, the evaluation results have shown the effectiveness of MAET and verified that MAET is a general LLM enhancement technology.
 LLM Tensors Proportion of $\mathcal{T}$ ES ZH BN TE Avg.
@@
 Table 8: The effect of only merging the specific LLM tensors during the transferring process (i.e., Eq.5) on multi-lingual mathematical tasks.
-Methods MMLU HumanEval MBPP OpenbookQA
-LLaMA-3 8B 60.85 35.98 46.60 65.00
-+ CPT 58.46 (-2.39) 28.66 (-7.32) 39.60 (-7.00) 61.40 (-3.60)
-+ MAET (Ours) 61.07 (+0.22) 35.98 (+0.00) 47.40 (+0.80) 65.00 (+0.00)
-Table 4: The out-of-domain performance comparison of different training methods to train LLaMA-3 8B on OpenWebMath.
-During the ability-enhancing process, previous methods will hurt the OOD abilities of LLM, while our MAET can maintain the corresponding abilities.
-Ability-related Sub-networks of LLM.
+Appendix F Ability-related Sub-networks of LLM
 To assess and probe the ability-related sub-networks of LLMs, we only transfer the specific tensors (i.e., tensors in self-attention and MLP mechanism) from the ability weight to the final models through Eq.
@@
 In the MLP layers of LLM, the gate mechanism (i.e., MLP Gate) will control the transmission of information and the down project mechanism (i.e., MLP Down) will integrate the knowledge from previous layers, so that transferring the MLP layers can achieve better performance on the downstream tasks.
-Out-of-Domain performance of MAET.
-We conduct experiments about adapting mathematical ability on the general LLM through MAET, and assess the performance on out-of-domain (OOD) tasks (i.e., MMLU (Hendrycks et al., 2021), HumanEval (Chen et al., 2021a), MBPP (Austin et al., 2021), and OpenbookQA (Mihaylov et al., 2018)), which can reflect and assess different abilities of LLMs.
-Results are presented in Table 4.
-We can observe that the performance of LLM on all evaluation tasks has decreased through the CPT training process, and the maximum decrease has been achieved 7.32% on the HumanEval task.
-One of the possible reasons is that LLaMA-3 has been trained on OpenWebMath during pre-training and the CPT process makes it overfit and forget the knowledge of other domains, hurting the performance on OOD tasks.
-In contrast, our proposed MAET achieves comparable and even better performance with backbone LLM in all downstream scenarios.
-Since we identify and update the key neurons related to the specific ability, the ability of LLM can be precisely enhanced, and this strategy also helps the OOD tasks needed for mathematical ability, e.g., mathematical sub-tasks in MMLU and code synthesis task MBPP.
-6 Conclusion
-In this paper, we presented MAET, which extracted the advanced ability-related weights from the LLM and supported simple addition and subtraction operations to transfer the ability across different languages.
-Concretely, MAET included two main stages, i.e., extraction and transfer.
-For the extraction stage, we located the key neurons and extracted the ability-related weights.
-Then, in the transfer stage, we identified the key parameter tensors and leveraged them to transfer the advanced ability into other LLMs.
-In this process, the multi-lingual ability-related training corpus is not required, and the experimental results have shown that our approach outperformed competitive baselines.
-As future work, we will consider better methods to identify the ability-related sub-network to decompose the abilities of LLMs and utilize an automated approach to determine the hyper-parameter.
-Besides, we will implement MAET on larger-scale models, and scenarios with more languages and requiring more abilities to evaluate its effectiveness.
-7 Limitations
-In this section, we discuss the limitations of our work.
-First, we only implement our approach MAET on 8B LLMs (i.e., LLaMA-3 8B), and do not adopt the LLMs with larger scales (e.g., 13B or 70B LLMs) in the experiment, due to the limitation of computational resources.
-We will test the effectiveness of our approach on these LLMs in the future.
-Second, we only evaluate our approach on two downstream tasks (i.e., mathematical and scientific reasoning tasks) in multi-lingual scenarios.
-Although they are challenging and widely-used testbeds, it is still meaningful to verify our methods on other tasks.
-Whereas, as we test the performance on diverse high-resource and low-resource languages, it can also provide comprehensive performance estimation for our approach in multi-lingual scenarios.
-Third, our approach is a general method for ability transferring across different domains, but in this work, we only consider the multi-lingual scenarios and obtain a multi-lingual LLM with the specific ability being enhanced.
-Forth, our approach also relies on continual pre-training the LLM.
-Although the training corpus is not very large, it also increases the cost.
-Fortunately, after we obtain the pre-trained weights, our following steps only need simple addition and subtraction operations for ability transferring, which is flexible for online deployment and application.
-In future work, we will focus on reducing the data requirement for the pre-training corpus and also improving the weights extracting efficiency.
-Finally, we do not consider the potential risk and ethics issues that might hurt the alignment of LLMs when using our approach.
-Actually, our approach is also applicable to transfer the alignment ability across languages.
-We will investigate to it in the future.
-Appendix A Empirical Study
-Figure 4: The results of empirical experiments.
-We present the loss of different training methods during the training process, the cosine similarity of LLM layers after being trained on Zhihu and Reddit, and the similarity of LLMs being trained on different training corpus.
-A surge of work (Zhang et al., 2024b; Xiao et al., 2024; Tang et al., 2024) has pointed out that LLMs sparsely activate the specific sub-modules to perform corresponding tasks.
-Based on these findings, we conduct empirical experiments to explore whether the specific sub-module, which is related to advanced abilities, can be extracted and transferred.
-We utilize the forum corpus (i.e., Zhihu for Chinese forum corpus and Reddit for English forum corpus) to continually pre-train LLMs, and then assess the training performance (i.e., the value of loss function) and similarity of LLM neurons.
-The forum corpus can be considered as containing the question-answering (QA) ability, which is necessary and important for LLMs.
-The results from Figure 4(a) have shown that only training the top 5% relevant neurons of LLMs can achieve the lower training loss and fit into the training set more quickly, indicating that LLMs contain the sub-module corresponding to the QA ability.
-Moreover, from Figure 4(b) and Figure 4(c), we can observe that the LLM trained on Zhihu has shown higher similarity with the LLM trained on Reddit than the LLM trained on Github (i.e., lower L1 Norm and higher cosine similarity), and the cosine similarity of different layers in LLM are largely different.
-According to the above results, we have found that the different sub-networks of LLMs control the different abilities, and precisely selecting the correct sub-module of LLMs will help the extraction of advanced abilities from the single-lingual corpus and the transfer of these abilities to multi-lingual scenarios.
-Concretely, although Zhihu and Reddit are in different languages, they will influence the similar sub-modules of LLM and make these sub-networks show high similarity with each other.
-These sub-networks can be referred to the ability-related sub-networks, which are slightly influenced by languages.
-Appendix B Details of Hyper-parameters
-We release all of the hyper-parameters during our experiment to better reproduce our proposed approach. Table 5 shows the details of hyper-parameters of different stages.
-Stage Hyper-Parameter Mathematical Tasks Scientific Tasks
-Extraction Learning Rate $5\times 10^{-5}$ $5\times 10^{-5}$
-Batch Size 1M Tokens 1M Tokens
-Training Steps 2B Tokens 2B Tokens
-$\alpha$ in Extraction 0.8 0.8
-$\beta$ in Extraction 0.2 0.2
-Ratio of Key Neurons 5% 5%
-Extraction Learning Rate $5\times 10^{-5}$ $5\times 10^{-5}$
-Batch Size 1M Tokens 1M Tokens
-Training Steps 2B Tokens 2B Tokens
-$\gamma$ in Transferring 0.2 0.2
-$\eta$ in Transferring 1.0 1.0
-Ratio of Key Neurons 80% 60%
-Table 5: The details of hyper-parameters in the evaluation.
-Appendix C Details of Dataset
-We present the statistical information of the datasets in Table 6.
-We mainly consider English, Spanish, Chinese, Bengali, and Telugu in our experiment, and utilized English as the in-domain language while others as the out-of-domain languages.
-For the evaluation datasets, we select MGSM and multi-lingual MMLU as the evaluation benchmarks, which contain the parallel data in different languages and are useful for multi-lingual complex tasks evaluation.
-Language Training Dataset (Tokens) Evaluation Dataset (Instances)
-General Corpus Ability-related Corpus Mathematical Tasks Scientific Tasks
-English 1.81B 1.30B (Math) / 1.82B (Sci) 250 1,245
-Spanish 1.81B - 250 1,232
-Chinese 1.80B - 250 1,229
-Bengali 1.81B - 250 1,137
-Telugu 1.81B - 250 1,036
-Table 6: The statistical information of the training and evaluation datasets.